OpenStudy (anonymous):

What is the average (arithmetic mean) of 8. 7. 7. 5, 3, 2 and 2? Please show answer step by step. Thank you. A. 3 4/7 B.4 5/6 C. 4 6/7 D. 5 E. 6 4/5

OpenStudy (anonymous):

Average is the sum of all your numbers divided by the total amount of numbers you have, so all your numbers added is 34, divide that by how many numbers you have which is 7, so 34/7... = 4.857 covert to fraction and your done

OpenStudy (anonymous):

How do I convert 4.857 to give me 4 6/7 as my answer?

OpenStudy (bahrom7893):

4857/1000 = 4 857/100

OpenStudy (bahrom7893):

wait im not sure about it, let me take it from 34/7

OpenStudy (bahrom7893):

34/7 = (28+6)/7 = (28/7) + (6/7) = 4 + (6/7) = 4 6/7

OpenStudy (anonymous):

Thanks!